During its first year of operations, Blossom Corporation had the following transactions pertaining to its common stock. Jan. 10 Issued 83,700 shares for cash at $7 per share. Mar. 1 Issued 5,000 shares to attorneys in payment of a bill for $37,000 for services rendered in helping the company to incorporate. July 1 Issued 32,100 shares for cash at $9 per share. Sept. 1 Issued 63,100 shares for cash at $11 per share. Prepare the journal entries for these transactions, assuming that the common stock has a par value of $5 per share. Prepare the journal entries for these transactions, assuming that the common stock is no-par with a stated value of $2 per share.
Answer #1
a
Debit Credit
Jan 10 Cash 585900 =83700*7
     Common Stock 418500 =83700*5
     Paid in Capital in excess of par value-Common Stock 167400
Mar 1 Organization Expense 37000
     Common Stock 25000 =5000*5
     Paid in Capital in excess of par value-Common Stock 12000
July 1 Cash 288900 =32100*9
     Common Stock 160500 =32100*5
     Paid in Capital in excess of par value-Common Stock 128400
Sept. 1 Cash 694100 =63100*11
     Common Stock 315500 =63100*5
     Paid in Capital in excess of par value-Common Stock 378600
b
Debit Credit
Jan 10 Cash 585900 =83700*7
     Common Stock 167400 =83700*2
     Paid in Capital in excess of stated value-Common Stock 418500
Mar 1 Organization Expense 37000
     Common Stock 10000 =5000*2
     Paid in Capital in excess of stated value-Common Stock 27000
July 1 Cash 288900 =32100*9
     Common Stock 64200 =32100*2
     Paid in Capital in excess of stated value-Common Stock 224700
Sept. 1 Cash 694100 =63100*11
     Common Stock 126200 =63100*2
     Paid in Capital in excess of stated value-Common Stock 567900